If a 75kg skydiver is falling and getting air resistance of 735N, what is the acceleration on the skydiver?
1
Expert's answer
2020-11-06T10:12:27-0500

According to Newton's second law, the acceleration is a consequence of the net force acting on a body, or


"a=\frac{F_\text{net}}{m}=\frac{mg-R}{m}=g-\frac Rm=0."

The skydiver is falling with terminal velocity (constant velocity, zero acceleration).